16 . Demand and Supply curves can be represented by:     Qd = 90-2P; and Qs = 3P.   Calculate the equilibrium price and quantity:

P = $18; and Q = 54
P = $90; and Q = 45
P = $30; and Q = 54
P = $30; and Q = 30
Answer #1

Demand Function

Qd = 90-2P

Supply Function

Qs = 3P  

Equilibrium condition is demand equals to supply.

Demand function = Supply function

90-2P = 3P

90 = 5P

P = 90/5 = $18

Quantity = 90 – 2 (18) = 54

Answer (a) P = $18; and Q = 54